OpenStudy (anonymous):

Which of these is the equation for a line tangent to the circle x^2 + y^2 = 20 at point (2, -4)? i got 2x+sq20 correct or no?

OpenStudy (freckles):

I would first find y' given x^2+y^2=20

OpenStudy (freckles):

then use the point (2,-4) to find the slope to the circle equation x^2+y^2=20 @ (2,-4) then use the point ... point slope form for a line is \[y-y_1=m(x-x_1) \\ \text{ we are given } (x_1,y_1)=(2,-4) \\ y-(-4)=m(x-2) \\ y+4=m(x-2) \\ y=m(x-2)-4 \] you need to find m

OpenStudy (freckles):

m is y' evaluated at (2,-4) (this is the slope at (2,-4))

OpenStudy (freckles):

can you find that?

OpenStudy (anonymous):

find the slope?

OpenStudy (anonymous):

or what do i find?

OpenStudy (freckles):

do you not know calculus?

OpenStudy (freckles):

If you don't we can do this algebraically

OpenStudy (anonymous):

its algebra class :p

OpenStudy (freckles):

ok y' is a calculus thing

OpenStudy (freckles):

so here we would have to look at this:  say this is our circle with center (0,0) and radius sqrt(20)

OpenStudy (freckles):

 a line that is drawn tangent to a circle is perpendicular to the radius I have drawn